# Regulating Sommerfeld resonances for multi-state systems and higher partial waves

### v1

@@ 1 Introduction
Ref. Blum:2016nrz demonstrated how to capture this effect and restore manifest unitarity for a single-state system with only $s$-wave annihilation, by modeling the effect of annihilation (and possibly other short-range physics) as a $\delta$-function potential with a complex coefficient, and solving the Schrödinger equation to obtain the wavefunction in the presence of this added term111 This calculation treats both long-range and short-range effects in a non-perturbative way. Related works can be found in the context of nuclear physics, including the analysis of $p\bar{p}$ bound states Trueman:1961zza ; Carbonell:1992wd , $p\bar{p}$ annihilation Carbonell:1993 ; Carbonell:1996vd ; Carbonell:1998ei ; Protasov:1999ei , and $pp$ scattering Kong:1998sx ; Kong:1999tw . . Refs. Braaten:2017gpq ; Braaten:2017kci ; Braaten:2017dwq developed a “zero range effective field theory” (ZREFT), allowing analytic study of Sommerfeld enhancement in the neighborhood of $s$-wave resonances, for DM that is a $SU(2)_{L}$ fermion triplet; the ZREFT parameterizes the behavior near the resonance in terms of a scattering length, and accounts for the annihilation effects naturally by allowing this scattering length to be complex. The ZREFT was restricted to $s$-wave resonances, but demonstrated the generalization of the unitarization prescription to a multi-state system. Ref. Chu:2019awd suggested a general parameterization for DM self-interaction cross sections in terms of the scattering length and effective range, and demonstrated how this prescription could be applied to the case of $s$-wave annihilation. Ref. Kamada:2023iol showed the correlation between the resonant self-scattering cross section and the Sommerfeld factor for the annihilation by using Watson’s theorem.
@@
The methodology of Ref. Blum:2016nrz did not generalize straightforwardly to higher partial waves, due to the lack of overlap between the $-\ell$ wavefunctions and a delta-function potential. In this work, we propose a new prescription for computing the corrected Sommerfeld enhancement, with conceptual similarities to the work of Refs. Agrawal:2020lea ; Parikh:2020ggm . The resulting expression is manifestly unitarity-preserving, requires no additional calculations beyond solving the Schrödinger equation for the long-range potential and knowledge of any relevant short-distance interactions, and can be applied to all partial waves and to multi-state systems.
@@
The central idea is that if there is a separation of scales between annihilation and the interactions mediating the long-range potential, then we can treat the annihilation (and possibly other short-distance physics) as being localized to a small-$r$ region $r<a$ (where $r$ is the separation between the interacting particles). Outside this region, we can solve the Schrödinger equation with the long-range potential as usual, and the only possible effect of any short-range interactions on this calculation will be to modify the boundary conditions at the matching radius. The modified boundary conditions can be related to the $S$-matrix element for scattering in the absence of the long-range potential, accounting only for the short-distance interactions, and assuming the short-distance interactions are weakly coupled, this short-distance $S$-matrix can be computed in the Born approximation (or equivalently, using perturbative quantum field theory (QFT)). A similar approach was used in Ref. Agrawal:2020lea to study DM elastic self-scattering in potentials with singular short-distance behavior.
@@
In Sec. 2 we introduce the method and apply it to the case where only a single state is relevant, obtain the regulated Sommerfeld enhancement for general $\ell$, demonstrate that we can recover the $\ell=0$ result of Ref. Blum:2016nrz , and work out some examples. In Sec. 3 we extend this approach to the case where multiple states are coupled by the long-range potential and can participate in the short-range interactions. We work out the example of wino dark matter in Sec. 4, and present our conclusions in Sec. 5. The appendices discuss the analytic properties of the basis solutions to the long-range Schrödinger equation; a modification of the variable phase method for numerically solving the Schrödinger equation for the case with modified short-distance boundary conditions; the approach we use for computing the short-distance scattering amplitude; some useful results on the partial-wave decomposition of the amplitude; some supplemental results for the wino example; and the generalization of our multi-state results to the case where the potential is Hermitian but not real.
@@ 2.4.1 Behavior of $u_{<,\ell}$, $F_{\ell}$, and $G_{\ell}$ at $r=a$
The Schrödinger equation at small $r$ provides recurrence relations among the coefficients of higher powers of $r$ as shown in Eqs. 172-174. Thus, the form of $F_{\ell}(r)$ (except for its overall prefactor) is uniquely determined by the Schrödinger equation at small $r$.
$G_{\ell}(r)$ is an irregular solution of the Schrödinger equation with long-range potential $V_{\rm long}(r)$. The recurrence relations for $G_{\ell}(r)$ are given in Eqs. 176-182. The coefficients of $r^{\ell}$, $r^{\ell-1}$, $\cdots$, $r^{-\ell+1}$ can be derived from those recurrence relations, and we can see that those term are subdominant at $r=a$, compared to the leading $r^{-\ell}$ term. On the other hand, it is impossible to get the relation between the coefficient of $r^{-\ell}$ and $r^{\ell+1}$ only from the recurrence relations near the origin. This is because the sum of an irregular solution and a regular solution is another irregular solution, satisfying the same Schrödinger equation. The coefficient of $r^{\ell}$ is determined by imposing the boundary condition at infinity (Eq. 17). For now, we just keep both the $r^{-\ell}$ term and $r^{\ell+1}$ term and write $G_{\ell}$ at $r=a$ as
@@ 2.4.2 Momentum scaling of terms in $k_{\ell}(p)$
Note that $c^{\prime}_{n}$ is an integer $\sim n!\log n$333 $c^{\prime}_{n}$ is defined as $\frac{d^{n}}{dx^{n}}(x^{n}\log x)=n!+c^{\prime}_{n}\log x$. The recurrence relation is $c^{\prime}_{1}=1$ and $c^{\prime}_{n+1}=(n+1)c^{\prime}_{n}+n!$. The solution is $c^{\prime}_{n}=n!(\gamma_{E}+\Psi(n))$ where $\gamma_{E}$ is Euler’s gamma. The difference between $\tilde{z}_{\ell}(p;a)$ and $z_{\ell}(p)+\tilde{c}_{\ell}x_{\ell}(p)$ is evaluated as
@@
In Eq. 38, $k_{\ell}(p_{0})$ and $\Delta z_{\ell}(p,p_{0};a)$ parametrize the effect of short-range physics and long-range physics, respectively. $\Delta z_{\ell}(p,p_{0};a)$ has a dependence on $a$ and so it seems that the separation between long-range physics and short-range physics is not complete. However, $\Delta z_{\ell}(p,p_{0};a)$ depends on $a$ via a term $[x_{\ell}(p)-x_{\ell}(p_{0})]\log(a/r_{0})$ and this term is always subdominant compared to the $k_{\ell}(p_{0})$ term. Therefore, we can safely ignore the $a$ dependence in $\Delta z_{\ell}(p,p_{0};a)$. Note that, in some special cases, $x_{\ell}(p)$ does not depend on momentum. For example, this happens for the case with $\ell=0$, or when the potential $V(r)$ has no $1/r$ behavior near the origin. In this case, the $a$-dependence completely disappears and the single parameter $k_{\ell}(p_{0})$ parameterizes the information of the short-range physics. However, for $\ell\geq 1$ and $V(r)$ containing a $1/r$ term, $x_{\ell}(p)$ in general does depend on $p$. For details, see App. A.3.
@@ 2.5 Comparison with literature
As mentioned previously, we can see that our formula in this limit is consistent with the conventional formula with Sommerfeld enhancement (e.g., Ref. ArkaniHamed:2008qn for the $s$-wave case and Refs. Iengo:2009ni ; Cassel:2009wt ; Slatyer:2009vg for $-\ell$ cases).
@@
Ref. Flores:2024sfy studies the preservation of unitarity in the presence of an effective imaginary potential [^4] A similar attempt at unitarization, to ensure consistency with the optical theorem, has been discussed in Ref. Kamada:2022zwb .. Whereas our method separates short-distance from long-distance physics, that work’s approach separates the real potential from the imaginary part, and does not require the imaginary part to be short-distance; in fact, the simplification employed in Eq. 60 of that work fails for absorptive contact interactions. Ref. Flores:2024sfy defines parameters $x_{\ell,\text{(un)reg}}=\sigma^{\text{elas}}_{\ell,\text{(un)reg}}/\sigma_{\ell}^{U}$, $y_{\ell,\text{(un)reg}}=\sigma^{\text{inel}}_{\ell,\text{(un)reg}}/\sigma_{\ell}^{U}$, where in our language $\sigma^{\text{inel}}_{\ell,\text{(un)reg}}$ is the inclusive (un)corrected annihilation cross section, $\sigma^{\text{elas}}_{\ell,\text{(un)reg}}$ is the (un)corrected elastic scattering cross section due to the long-range potential, and $\sigma_{\ell}^{U}=2^{\delta}4\pi(2\ell+1)/p^{2}$ is the partial wave unitarity bound, where $\delta$ is 0 or 1 depending on whether the incoming state contains distinguishable or identical particles (so $c_{i}=2^{\delta}$ in our notation). Our Eqs. 24, 25 thus predict:
@@
These results are identical to the expressions given in Eq. 60 of Ref. Flores:2024sfy . In general our $k_{\ell}^{-1}(p)$ will not be purely imaginary and is not governed purely by the short-range physics, receiving irreducible contributions from the long-range real potential; however, it may often be true that the term arising from the short-range physics is the dominant one.
@@ 2.6.1 Spherical well potential
Now let us discuss the $\tilde{z}_{\ell}$ function. Since the potential $V(r)$ around $r=0$ is flat and there is no $1/r$ term, $x_{\ell}$ is equal to 00 in this case. Thus, $\tilde{z}_{\ell}$ is finite in the limit $a\to 0$. By using Eq. 35, we obtain
@@
We can see that the factor $\left|1+\tilde{z}_{\ell}(p;0)/k_{\ell}(p_{0})\right|^{-2}$ cancels the divergent behavior of $C_{\ell}$ at low velocity, and $\sigma_{{\rm ann},\ell}/\sigma_{{\rm ann},\ell}^{(0)}$ becomes a momentum-independent constant in the limit of $p\to 0$. In Fig. 4, we show the Sommerfeld enhancement computed by the conventional formula Eq. 45, the full formula Eq. 42, and the simplified formula Eqs. 68, 69. We can see that the divergent behavior in the conventional formula (which can lead to unitarity violation) is regularized in the full formula given in Eq. 45. For the $s$-wave case, the behavior of the full formula is well captured by the simplified formula in Eq. 68. For higher partial waves, the simplified formula in Eq. 69 describes the behavior of the full formula fairly well except for the points where $\tilde{z}_{\ell}$ crosses 00 and flips sign.
@@ 2.6.2 Coulomb potential
$x_{\ell}(p)=-2p^{\ell-1}a_{B}^{-\ell-2}g_{\ell}\left(-\frac{1}{pa_{B}}\right).$ (76)
@@
$\tilde{z}_{0}=\frac{p^{0}C_{0}}{2^{0}0!}\frac{d}{dr}G_{0}\biggr{|}_{r=a}=$ $pa_{B}x_{0}(p)\left[\log\left(\frac{2a}{a_{B}}\right)+2\gamma_{E}+h\left(\frac{1}{pa_{B}}\right)\right],$ (77)
$\tilde{z}_{1}=\frac{p^{1}C_{1}}{2^{1}1!}\frac{d^{3}}{dr^{3}}r^{1}G_{1}\biggr{|}_{r=a}=$ $p^{2}a_{B}^{2}x_{1}(p)\left[\log\left(\frac{2a}{a_{B}}\right)+2\gamma_{E}+h\left(\frac{1}{pa_{B}}\right)\right]$
$-\frac{p^{2}}{2a_{B}},$ (78)
$\tilde{z}_{2}=\frac{p^{2}C_{2}}{2^{2}2!}\frac{d^{5}}{dr^{5}}r^{2}G_{2}\biggr{|}_{r=a}=$ $p^{3}a_{B}^{3}x_{2}(p)\left[\log\left(\frac{2a}{a_{B}}\right)+2\gamma_{E}+h\left(\frac{1}{pa_{B}}\right)\right]$
$-\frac{p^{2}}{16a_{B}^{3}}(4+13p^{2}a_{B}^{2}),$ (79)
$\tilde{z}_{3}=\frac{p^{3}C_{3}}{2^{3}3!}\frac{d^{7}}{dr^{7}}r^{3}G_{3}\biggr{|}_{r=a}=$ $p^{4}a_{B}^{4}x_{3}(p)\left[\log\left(\frac{2a}{a_{B}}\right)+2\gamma_{E}+h\left(\frac{1}{pa_{B}}\right)\right]$
$-\frac{p^{2}}{48a_{B}^{5}}(2+23p^{2}a_{B}^{2}+50p^{4}a_{B}^{4}),$ (80)
@@
Here $h(\eta)$ is defined as below Eq. 23. Eq. 77 shows that the $\log a$ term in $G_{\ell}$ has a momentum-independent coefficient for $\ell=0$. On the other hand, Eqs. 7880 show the coefficient is momentum-dependent for $\ell\geq 1$ (although at least in this example, the coefficient becomes momentum-independent for $p\ll 1/a_{B}$ for general $\ell$, as $g_{\ell}(\eta)\propto\eta^{2\ell}$ for large $\eta$). Thus we cannot take the $a\to 0$ limit (and absorb the log term into the high-momentum amplitude) naively for $\ell\geq 1$ as discussed in Sec. 2.4. $\tilde{z}_{\ell}$ is constructed from a polynomial in $p$ and the function $h(1/pa_{B})$. Note that $h(\eta)$ does not have a singularity for $\eta>0$ and its asymptotic behavior is $h(\eta)\simeq-\log\eta-\gamma_{E}$ for $\eta\ll 1$ and $h(\eta)\simeq 1/12\eta^{2}$ for $\eta\gg 1$. Thus, $\tilde{z}_{\ell}$ also does not have divergent behavior at small $p$. By using dimensional analysis, $|\tilde{z}_{\ell}|$ should be bounded by $[{\rm max}(p,a_{B}^{-1})]^{2\ell+1}$. Thus, as long as the short-range effect $k_{\ell}^{-1}$ is perturbative, $|k_{\ell}|\gg|z_{\ell}|$ should be satisfied and the effect of $z_{\ell}$ is negligible. The numerical value of $\tilde{z}_{\ell}$ is shown in Fig. 5, which shows that a rough behavior of $\tilde{z}_{\ell}$ is given as

### v3

@@ 1 Introduction
Ref. Blum:2016nrz demonstrated how to capture this effect and restore manifest unitarity for a single-state system with only $s$-wave annihilation, by modeling the effect of annihilation (and possibly other short-range physics) as a $\delta$-function potential with a complex coefficient, and solving the Schrödinger equation to obtain the wavefunction in the presence of this added term [^1] This calculation treats both long-range and short-range effects in a non-perturbative way. Related works can be found in the context of nuclear physics, including the analysis of $p\bar{p}$ bound states Trueman:1961zza ; Carbonell:1992wd , $p\bar{p}$ annihilation Carbonell:1993 ; Carbonell:1996vd ; Carbonell:1998ei ; Protasov:1999ei , and $pp$ scattering Kong:1998sx ; Kong:1999tw . . Refs. Braaten:2017gpq ; Braaten:2017kci ; Braaten:2017dwq developed a “zero range effective field theory” (ZREFT), allowing analytic study of Sommerfeld enhancement in the neighborhood of $s$-wave resonances, for DM that is a $SU(2)_{L}$ fermion triplet; the ZREFT parameterizes the behavior near the resonance in terms of a scattering length, and accounts for the annihilation effects naturally by allowing this scattering length to be complex. The ZREFT was restricted to $s$-wave resonances, but demonstrated the generalization of the unitarization prescription to a multi-state system. Ref. Chu:2019awd suggested a general parameterization for DM self-interaction cross sections in terms of the scattering length and effective range, and demonstrated how this prescription could be applied to the case of $s$-wave annihilation. Ref. Kamada:2023iol showed the correlation between the resonant self-scattering cross section and the Sommerfeld factor for the annihilation by using Watson’s theorem.
@@
The methodology of Ref. Blum:2016nrz did not generalize straightforwardly to higher partial waves, due to the lack of overlap between the higher-$\ell$ wavefunctions and a delta-function potential. In this work, we propose a new prescription for computing the corrected Sommerfeld enhancement, with conceptual similarities to the work of Refs. Agrawal:2020lea ; Parikh:2020ggm . The resulting expression is manifestly unitarity-preserving, requires no additional calculations beyond solving the Schrödinger equation for the long-range potential and knowledge of any relevant short-distance interactions, and can be applied to all partial waves and to multi-state systems. Our work complements that of Ref. Flores:2024sfy , which showed how to extend the single-state unitarization calculation to higher partial waves via an approach based on the generalized optical theorem.
@@
The central idea of our work is that if there is a separation of scales between annihilation and the interactions mediating the long-range potential, then we can treat the annihilation (and possibly other short-distance physics) as being localized to a small-$r$ region $r<a$ (where $r$ is the separation between the interacting particles). Outside this region, we can solve the Schrödinger equation with the long-range potential as usual, and the only possible effect of any short-range interactions on this calculation will be to modify the boundary conditions at the matching radius. The modified boundary conditions can be related to the $S$-matrix element for scattering in the absence of the long-range potential, accounting only for the short-distance interactions, and assuming the short-distance interactions are weakly coupled, this short-distance $S$-matrix can be computed in the Born approximation (or equivalently, using perturbative quantum field theory (QFT)). A similar approach was used in Ref. Agrawal:2020lea to study DM elastic self-scattering in potentials with singular short-distance behavior.
@@
In Sec. 2 we introduce the method and apply it to the case where only a single state is relevant, obtain the regulated Sommerfeld enhancement for general $\ell$, demonstrate that we can recover the $\ell=0$ result of Ref. Blum:2016nrz , and work out some examples. In this section we also develop an analytic understanding of the momentum dependence of the enhancement in the low-momentum limit. In Sec. 3 we extend this approach to the case where multiple states are coupled by the long-range potential and can participate in the short-range interactions, and show how to write the full regulated annihilation cross section in terms of the matrix elements for short-distance interactions and factors derived from the long-range Schrödinger equation. We work out the example of wino dark matter in Sec. 4, and present our conclusions in Sec. 5.
Appendices A and B collect results that we use pertaining to solutions of the long-range Schrödinger equation; App. A discusses the analytic properties of the basis solutions to the long-range Schrödinger equation, while App. B presents a modification of the variable phase method for numerically solving the Schrödinger equation for the case with modified short-distance boundary conditions. App. C works out in detail the comparison between our single-state results and those of Ref. Flores:2024sfy , and shows that under a certain set of assumptions and approximations they are equivalent. Apps. D and E discuss, respectively, the approach we use for computing the short-distance scattering amplitude and matching between the QFT and quantum mechanics calculations, and the interpretation of the short-distance amplitude when we use a convention where the long-range potential still has some support at $r<a$. We present some useful results on the partial-wave decomposition of the amplitude in App.  F; some supplemental results for the wino example in App. G; and the generalization of our multi-state results to the case where the potential is Hermitian but not real in App. H.
@@ 2.4.1 Behavior of $u_{<,\ell}$, $F_{\ell}$, and $G_{\ell}$ at $r=a$
The Schrödinger equation at small $r$ provides recurrence relations among the coefficients of higher powers of $r$ as shown in Eqs. 185-187. Thus, the form of $F_{\ell}(r)$ (except for its overall prefactor) is uniquely determined by the Schrödinger equation at small $r$.
$G_{\ell}(r)$ is an irregular solution of the Schrödinger equation with long-range potential $V_{\rm long}(r)$. The recurrence relations for $G_{\ell}(r)$ are given in Eqs. 189-195. The coefficients of $r^{\ell}$, $r^{\ell-1}$, $\cdots$, $r^{-\ell+1}$ can be derived from those recurrence relations, and we can see that those terms are subdominant at $r=a$, compared to the leading $r^{-\ell}$ term. In more detail, we expect the coefficients $g_{\ell,k}$ (which are normalized so that the coefficient for the $r^{-\ell}$ term is dimensionless and $\mathcal{O}(1)$, see Eq. 189) to take the form of polynomials in the momentum $p$ and the coefficients $v_{k}$ defined from the expansion of the potential in Eq. 183. These coefficients do not have any singular behavior near the resonance points, and have positive mass dimension $k+2$. If we write $v_{k}\sim p_{k}^{k+2}$, where $p_{k}$ is a momentum scale associated with the potential, we expect $p_{k}\ll\mu\sim 1/a$ by our assumption that the potential is long-range (note that for the Coulomb potential, $v_{-1}$ is just set by the inverse Bohr radius). We thus expect the coefficients to be built from positive powers of these $p_{k}$ factors and the small momentum $p$, and the resulting terms to be suppressed (relative to the $r^{-\ell}$ term) for $r\sim a$ by products of factors of the form $p_{k}a\ll 1$, $pa\ll 1$.
However, this argument breaks down for the $r^{\ell+1}$ term, because it is impossible to get the relation between the coefficient of $r^{-\ell}$ and $r^{\ell+1}$ only from the recurrence relations near the origin. This is because the sum of an irregular solution and a regular solution is another irregular solution, satisfying the same Schrödinger equation. This means that the $r^{\ell+1}$ term can in principle have a coefficient that is parametrically enhanced, and a value much larger than one would naively expect from dimensional analysis + the assumption that the only relevant momentum scales are small compared to $1/a$.
The coefficient of $r^{\ell+1}$ is determined by imposing the boundary condition at infinity (Eq. 17). Thus, although $r^{\ell}$, $r^{\ell-1}$, $\cdots$, $r^{-\ell+1}$ are subdominant at $r=a$, we cannot conclude whether the $r^{\ell+1}$ term is subdominant or not. A general expansion of $G_{\ell}$ is given in Eq. 188. For now, we just keep both the $r^{-\ell}$ term and $r^{\ell+1}$ term and write $G_{\ell}$ at $r=a$ as
@@ 2.4.2 Momentum scaling of terms in $k_{\ell}(p)$
[…]
@@
In Eq. 38, $k_{\ell}(p_{0})$ and $\Delta z_{\ell}(p,p_{0};a)$ parametrize the effect of short-range physics and long-range physics, respectively. $\Delta z_{\ell}(p,p_{0};a)$ has a dependence on $a$ and so it seems that the separation between long-range physics and short-range physics is not complete. However, $\Delta z_{\ell}(p,p_{0};a)$ depends on $a$ via a term $x_{\ell}(p)-x_{\ell}(p_{0})]\log(a/r_{0})$ and this term is always subdominant compared to the $k_{\ell}(p_{0})$ term. Therefore, we can safely ignore the $a$ dependence in $\Delta z_{\ell}(p,p_{0};a)$. Note that, in some special cases, $x_{\ell}(p)$ does not depend on momentum. For example, this happens for the case with $\ell=0$, or when the potential $V(r)$ has no $1/r$ behavior near the origin. In this case, the $a$-dependence completely disappears and the single parameter $k_{\ell}(p_{0})$ parameterizes the information of the short-range physics. However, for $\ell\geq 1$ and $V(r)$ containing a $1/r$ term, $x_{\ell}(p)$ in general does depend on $p$. For details, see App. A.3.
@@ 2.5 Comparison with literature
As mentioned previously, we can see that our formula in this limit is consistent with the conventional formula with Sommerfeld enhancement (e.g., Ref. ArkaniHamed:2008qn for the $s$-wave case and Refs. Iengo:2009ni ; Cassel:2009wt ; Slatyer:2009vg for higher-$\ell$ cases).
@@
Ref. Flores:2024sfy studies the preservation of unitarity in the presence of an effective imaginary potential [^4] A similar attempt at unitarization, to ensure consistency with the optical theorem, has been discussed in Ref. Kamada:2022zwb .. Whereas our method separates short-distance from long-distance physics, that work’s approach separates the real potential from the imaginary part, and does not require the imaginary part to be short-distance; in fact, the assumption needed in that work to obtain a simple expression for the corrected cross section fails for absorptive contact interactions. Ref. Flores:2024sfy defines parameters $x_{\ell,\text{(un)reg}}=\sigma^{\text{elas}}_{\ell,\text{(un)reg}}/\sigma_{\ell}^{U}$, $y_{\ell,\text{(un)reg}}=\sigma^{\text{inel}}_{\ell,\text{(un)reg}}/\sigma_{\ell}^{U}$, where in our language $\sigma^{\text{inel}}_{\ell,\text{(un)reg}}$ is the inclusive (un)corrected annihilation cross section, $\sigma^{\text{elas}}_{\ell,\text{(un)reg}}$ is the (un)corrected elastic scattering cross section due to the long-range potential, and $\sigma_{\ell}^{U}=2^{\delta}4\pi(2\ell+1)/p^{2}$ is the partial wave unitarity bound, where $\delta$ is 0 or 1 depending on whether the incoming state contains distinguishable or identical particles (so $c_{i}=2^{\delta}$ in our notation). Our Eqs. 24, 25 thus predict:
@@
These results are identical to the expressions given in Eq. 60 of Ref. Flores:2024sfy . For a detailed comparison of our wave function $u_{\ell}(r)$ and Ref. Flores:2024sfy , see App. C. In general our $k_{\ell}^{-1}(p)$ will not be purely imaginary and is not governed purely by the short-range physics, receiving irreducible contributions from the long-range real potential; however, it may often be true that the term arising from the short-range physics is the dominant one.
@@ 2.6.1 Spherical well potential
Now let us discuss the $\tilde{z}_{\ell}$ function. Since the potential $V(r)$ around $r=0$ is flat and there is no $1/r$ term, $x_{\ell}$ is equal to 0 in this case. Thus, $\tilde{z}_{\ell}$ is finite in the limit $a\to 0$. By using Eq. 35, we obtain
@@
We can see that the factor $\left|1+\tilde{z}_{\ell}(p;0)/k_{\ell}(p_{0})\right|^{-2}$ cancels the divergent behavior of $C_{\ell}$ at low velocity, and $\sigma_{{\rm ann},\ell}/\sigma_{{\rm ann},\ell}^{(0)}$ becomes a momentum-independent constant in the limit of $p\to 0$. In Fig. 4, we show the Sommerfeld enhancement computed by the conventional formula Eq. 45, the full formula Eq. 42, and the simplified formula Eqs. 68, 69. We can see that the divergent behavior in the conventional formula (which can lead to unitarity violation) is regularized in the full formula given in Eq. 45. For the $s$-wave case, the behavior of the full formula is well captured by the simplified formula in Eq. 68. For higher partial waves, the simplified formula in Eq. 69 describes the behavior of the full formula fairly well except for the points where $\tilde{z}_{\ell}$ crosses 0 and flips sign.
@@ 2.6.2 Coulomb potential
$x_{\ell}(p)=-2p^{2\ell}a_{B}^{-1}g_{\ell}\left(-\frac{1}{pa_{B}}\right).$ (76)
@@
$\tilde{z}_{0}=\frac{p^{0}C_{0}}{2^{0}0!}\frac{d}{dr}G_{0}\biggr|_{r=a}=$ $x_{0}(p)\left[\log\left(\frac{2a}{a_{B}}\right)+2\gamma_{E}+h\left(\frac{1}{pa_{B}}\right)\right],$ (77)
$\tilde{z}_{1}=\frac{p^{1}C_{1}}{2^{1}1!}\frac{d^{3}}{dr^{3}}r^{1}G_{1}\biggr|_{r=a}=$ $x_{1}(p)\left[\log\left(\frac{2a}{a_{B}}\right)+2\gamma_{E}+h\left(\frac{1}{pa_{B}}\right)\right]-\frac{p^{2}}{2a_{B}},$ (78)
$\tilde{z}_{2}=\frac{p^{2}C_{2}}{2^{2}2!}\frac{d^{5}}{dr^{5}}r^{2}G_{2}\biggr|_{r=a}=$ $x_{2}(p)\left[\log\left(\frac{2a}{a_{B}}\right)+2\gamma_{E}+h\left(\frac{1}{pa_{B}}\right)\right]-\frac{p^{2}}{16a_{B}^{3}}(4+13p^{2}a_{B}^{2}),$ (79)
$\tilde{z}_{3}=\frac{p^{3}C_{3}}{2^{3}3!}\frac{d^{7}}{dr^{7}}r^{3}G_{3}\biggr|_{r=a}=$ $x_{3}(p)\left[\log\left(\frac{2a}{a_{B}}\right)+2\gamma_{E}+h\left(\frac{1}{pa_{B}}\right)\right]$
$\qquad-\frac{p^{2}}{48a_{B}^{5}}(2+23p^{2}a_{B}^{2}+50p^{4}a_{B}^{4}),$ (80)
@@
[…]